OpenStudy (anonymous):

need to solve... y=x^2-5x+10 and y=-x+10

OpenStudy (amistre64):

since y=y, set them equal to each other and play the quadratic

OpenStudy (amistre64):

its even a simple quadratic .... since y = a+10 and y=b+10, then it stands to reason that a=b

OpenStudy (amistre64):

x^2-5x = -x